OpenStudy (alyssajobug):

Solve. 4y - 9 - 6y = 2(y + 5) - 3 A. y = -3 B. y = 2 C. y = -4 D. y = 5

OpenStudy (anonymous):

4y - 9 -6y = 2(y+5)-3 Distribute and Combine Like Terms -2y - 9 = 2y +7 Move the y's so that they're on the same side of the = sign -4y = 16 Dived by -4 y = -4 Answer: C, y = -4